OpenStudy (anonymous):

Which of the following is a solution of 9x^2-5x = -3? (Quadratic formula)

hartnn (hartnn):

\(9x^2-5x +3=0 \) Compare this quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

wait that didn't help??

OpenStudy (anonymous):

i obviously know that

OpenStudy (anonymous):

but the answers have "i" in them so I'm confused

OpenStudy (anonymous):

@hartnn there's no point in just typing the quadratic formula. if i know i have to use it for the problem, i can just look it up. if ur not gonna help, then..

OpenStudy (anonymous):

I mean, like, i just don't need info I already have

hartnn (hartnn):

yes, first tell me what you get for b^2-4ac ?

OpenStudy (anonymous):

the choices are up there^^

hartnn (hartnn):

could you find a,b,c ?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

yeah ok, so a=9, b= -5, c= -3

OpenStudy (anonymous):

and what do you mean "the two roots of x"

OpenStudy (anonymous):

why does it say x 1,2 in front of the formula

hartnn (hartnn):

that there are 2 solutions for x

hartnn (hartnn):

c is just +3

OpenStudy (anonymous):

ohh, i forgot to do the discriminant

OpenStudy (anonymous):

it is? why

OpenStudy (anonymous):

ohh i get it

OpenStudy (anonymous):

wait but i got a negative discriminant which means there are no real solutions...

OpenStudy (anonymous):

how did u get a positive one

hartnn (hartnn):

for this, there are actually no real solutions, thats why there is "i" in the answers which is imaginary number

hartnn (hartnn):

so, what you got for b^2-4ac ?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

i got -98

OpenStudy (anonymous):

and negative means two imaginary solutions

hartnn (hartnn):

try again ?

OpenStudy (anonymous):

i did it wrong? ok

hartnn (hartnn):

25 -4*3*9 = ... ?

OpenStudy (anonymous):

ohhh i see my mistake hold on

OpenStudy (anonymous):

wait ... its 567

hartnn (hartnn):

nopes, ist -83

OpenStudy (anonymous):

ugh i didnt do order of operations

OpenStudy (anonymous):

ok lets keep going

hartnn (hartnn):

so, you know \(i =\sqrt {-1}\) ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh so now we can look at the choices

hartnn (hartnn):

so, what about \(\sqrt{-83}\) ? yes...

OpenStudy (anonymous):

so sqrt -1 times -83

OpenStudy (anonymous):

or is it 83

OpenStudy (anonymous):

i think its b

OpenStudy (anonymous):

bcuz the rest have a positive 5

OpenStudy (anonymous):

and for the b value it is negative 5

hartnn (hartnn):

-b means - (-5) = 5 so the answer should have +5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

but before we said it was negative

OpenStudy (anonymous):

can u like write it out??

OpenStudy (anonymous):

where did the negative sign before the (-5) come from

hartnn (hartnn):

yes

hartnn (hartnn):

\(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}} \\ a=9, b=-5,c =3 \\ \huge{x_{1,2}=\frac{-(-5) \pm i\sqrt{83}}{2*9}}=...?\)

OpenStudy (anonymous):

but why -(-5)?

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

im so stupid :((

OpenStudy (anonymous):

so i sqrt 83 will be -1 times 83 right

hartnn (hartnn):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

will 83 be under a sq rt

OpenStudy (anonymous):